# Random Access for LEO Satellite Communication Systems via Deep Learning

### v1

@@ IV-A Preamble Transmission and Reception
where $P_{d}$ is the transmit power, and $\hat{\tau}_{d}$ and $\hat{f}_{d}$ denote the estimated timing advance and Doppler shift, respectively. $T_{\mathrm{s}}$ denotes the sampling period. We consider a multi-path channel model consisting of $L_{d}$ paths for the $d$th user. The SBS is equipped with $N_{\mathrm{ant}}$ antennas, each of which receives the superposition of signals reflected through multiple paths. Each path is characterized by a channel coefficient $h_{d,\ell}^{j}$ and a propagation delay $\tau_{d,\ell}^{j}$ at the $j$th antenna. The signal received at antenna $j$ of the SBS is modeled as
@@
where $D$ is the number of active users, $\tau_{d}^{\mathrm{e}}$ and $f_{d}^{\mathrm{e}}$ represent the residual errors in timing and frequency pre-compensation, and $w[n]$ is additive white Gaussian noise with zero mean and variance $\sigma_{w}^{2}$. Following the conventional detection procedure, the SBS computes the correlation between the received signal and each ZC sequence. The correlation at lag $m$ for root index $r$ at antenna $j$ is given by
@@ IV-B Preamble Collision Classifier
As mentioned, collisions are not detected by traditional schemes until Step 3. To detect preamble collisions at an early stage, we design a deep learning-based classifier that analyzes the correlation values in (8). For a given root index and cyclic shift, the classifier considers a total $(K+1)$ classes, where each class corresponds to the following scenarios:
@@
  • Class 1: A single user has used the preamble, indicating successful use without collision.
@@
  • Class 2 to $\boldsymbol{(K-1)}$: The preamble has been used by 2 to (K1)(K-1) users, indicting a collision has occurred.
@@
  • Class $\boldsymbol{K}$: The preamble has been used by $K$ or more users, also indicating a collision has occurred.
@@
The advantage of post-processing the correlation output, as opposed to relying solely on traditional threshold-based detection, is that it not only identifies whether a collision has occurred but also determines the number of colliding users.
@@
In realizing such a classifier, our goal is to propose a framework that can be easily integrated into legacy random access procedures by maintaining the core signaling of the conventional random access framework. Consequently, the input data for the classifier should be correlation values (8). Since each user independently selects a preamble and collisions across different preambles are independent events, we use a single ZCZ as the input data. A key point to note, as previously mentioned, is that in SatCom scenarios, timing advance is pre-compensated by the users during the preamble transmission stage. As shown in (5), the ZCZ size in SatCom scenarios is expected to be considerably smaller than in terrestrial networks, primarily because the otherwise large RTT uncertainty is mitigated through timing advance pre-compensation. For example, while the typical ZCZ size in terrestrial networks is approximately 24 samples [TWC21_DLRA], it is assumed to be only 8 samples in SatCom scenarios [TVT23_TwoStage, TVT25_MTeCD]. This limited ZCZ size imposes a fundamental constraint on the amount of information that can be extracted for preamble collision detection. More specifically, the smaller number of samples reduces the temporal resolution available to distinguish multiple colliding signals. As a result, subtle differences in time alignment and energy distribution between users become harder to capture, making it more challenging to detect collisions based solely on the correlation values.
@@
Fortunately, LEO satellites employ multi-antenna systems [COMST23_LEOMIMO], which provide multiple observations of the transmitted preambles. In terrestrial networks, the correlation values obtained from multiple antennas or ports are typically averaged for threshold-based detection [Book_RABO]. Similarly, existing deep learning-based schemes have also used the averaged correlation values as input data [TWC21_DLRA, TWC23_DLDouble]. Averaging enables receive antenna diversity but at the cost of reduced input dimensionality, which hinders full utilization of the original signal information. To fully exploit the information contained within preamble signals received across multiple antennas, we use the correlation values from each antenna as input data for the classifier. This preserves the spatial and temporal diversity in the received signals, providing the classifier with richer information for more accurate collision detection.
@@
However, this also increases the input data dimensionality, which in turn raises the network complexity. To mitigate this, we employ a 1D convolutional layer for extracting features of the preamble collision pattern. Compared to fully connected layers, 1D convolutional layers require significantly fewer parameters, which allows the network to maintain lower complexity despite the increased input size. Moreover, convolutional layers are well-suited for this task, as they can effectively extract local features which are useful in classification. In fading channels, for example, the received preamble signal spreads around its peak due to multipath effects, making it critical to capture the local features of the received signal. In addition, the kernel size of the convolutional layer can be matched to the typical delay spread observed in the SatCom channel, allowing the network to focus on the most relevant temporal dependencies for collision detection. This makes the 1D convolutional structure particularly effective for capturing the subtle signal differences that arise in the limited ZCZ region.
The structure of our proposed early preamble collision classifier is shown in Fig. 4. The input data consists of the correlation values of the preamble’s ZCZ for each antenna, denoted as $c_{r}^{j}\mathopen{}\mathclose{{\left[m}}\right]$, where $m\in\mathopen{}\mathclose{{\left[\mathopen{}\mathclose{{\left(i-1}}\right)N_{\mathrm{CS}},iN_{\mathrm{CS}}-1}}\right]$. The input layer is of size $N_{\mathrm{ant}}\times N_{\mathrm{CS}}$, where $N_{\mathrm{CS}}$ is set to 8 according to the condition in (5[TVT23_TwoStage]. The kernel size of the 1D convolutional layer is set to 3, which reflects the typical delay spread of the SatCom channel as reported in [38.811] and captures local dependencies across adjacent samples. After the input layer, the data passes through a 1D convolutional layer with a kernel size of 3, stride 1, padding 1, and 16 channels. The rectified linear unit (ReLU), defined as $\mathrm{ReLU}(x)=\max\mathopen{}\mathclose{{\left(0,x}}\right)$, is used as the activation function. The next layer is another 1D convolutional layer with the same kernel size, stride, and padding, but with 32 channels, designed to extract high-level features of the preamble collision pattern. After passing through the activation function, the data is fed into a fully connected layer for classification. The output size of the fully connected layer is $K+1$, corresponding to the number of classes. The final output values $z_{k}$ are processed using the softmax function $\sigma(z_{k})\triangleq\frac{e^{z_{k}}}{\sum^{K}_{i=0}e^{z_{i}}}$ to compute the probability of each class. Cross-entropy is used for the loss function, i.e.,
@@ IV-C Opportunistic Transmission Scheme for Step 3
    The SBS transmits the RAR message to the users, including the preamble index, PUSCH grant, a temporal identifier linking the RAR to the original preamble, and the calculated Step 3 transmission probability for each preamble index.
@@
    Each user refers to the RAR message corresponding to its transmitted preamble index and decides whether to attempt Step 3 or perform random backoff based on the provided transmission probability.
Based on the estimated number of collided users, we now derive the Step 3 transmission probability that maximizes the random access success probability. This probability allows the system to balance the trade-off between minimizing unnecessary Step 3 transmissions and maintaining a high success probability, even in the presence of collisions. Let the number of collided users in the ZCZ of a given preamble be classified as $\hat{k}$, and let the Step 3 transmission probability for this preamble be denoted as $P$. For notational simplicity, we omit the preamble index in the following derivation, without loss of generality. We define successful random access as the event in which exactly one of the users who selected the same preamble attempts Step 3, while all others perform backoff. Based on this definition, the random access success probability for the preamble is given as
@@ V Simulation Results
This section evaluates the performance of the proposed deep learning-based random access framework for LEO satellite communication systems. The tapped delay line (TDL) channel model recommended for SatCom link-level evaluation [38.811] is employed as the channel model. We assume that GNSS-based timing advance and frequency offset pre-compensation are performed at the user side [38.811]. Other random access-related parameters are summarized in Table I.
@@
TABLE I: Simulation parameter settings [TVT23_TwoStage, TWC17_NORA, 36.213, IoTJ21_ERA, 38.821].
@@
The dataset for training the collision classifier consists of ZCZ correlation values generated under the TDL-B and TDL-D channel models, which are representative of NLoS and LoS-dominant propagation environments, respectively. Data are collected for all classes from 0 to $K$, corresponding to $K+1$ collision scenarios. To accurately simulate the reception of preambles under these scenarios, appropriate transmit power and noise configurations are required. In 3GPP-compliant systems, preamble transmission is subject to open-loop power control, where the transmit power is adjusted to satisfy specific detection criteria. To capture this in our simulation, we configure the transmit power and noise level based on 3GPP open-loop power control and detection requirements. Specifically, the noise level is set to satisfy the 3GPP-specified performance targets of misdetection probability below 1% and false alarm probability below 0.1% [Book_RABO]. Fig. 5 shows the misdetection and false alarm probabilities of the proposed classifier under various noise levels, where the misdetection probability is defined as the likelihood of classifying a non-zero class as class 0, and the false alarm probability is the likelihood of classifying a true class 0 as another class. To satisfy 3GPP requirements [Book_RABO], the minimum required signal-to-noise ratio (SNR) is found to be at least 13-13 dB. Accordingly, 100,000 data samples were collected at each SNRs of {13-13, 12-12, 11-11, 10-10} dB, of which 70% were used for training and 30% for testing. Additionally, as in [TWC21_DLRA], to demonstrate the generalization performance of the trained classifier, 30,000 samples were collected at each SNRs of {9-9, 8-8, 7-7, 6-6, 5-5, 4-4} dB as a separate test set. Hyperparameters including learning rate, batch size, number of epochs, and Adam optimizer settings were tuned using $K$-fold cross-validation. The final values were set to a learning rate of $10^{-3}$, a batch size of 32, 20 epochs, and Adam optimizer parameters $(\beta_{1},\beta_{2})=(0.9,0.999)$, respectively.
@@
  • Deep learning-based random access framework for terrestrial networks (DRA) [TWC21_DLRA]: This scheme employs a deep learning-based classifier for collision detection and uses a separate neural network to extract timing advance information, which is then used to identify users and allocate distinct Step 3 grants. However, it assumes that timing advance is stationary and known to users in advance, making it unsuitable for LEO SatCom systems. Therefore, it is only used to compare preamble collision classification accuracy.
  • Deep learning-based double contention random access (DCRA) [TWC23_DLDouble]: This scheme extends DRA [TWC21_DLRA] by adopting a wider and deeper neural network for improved collision classification. When two users are classified as collided, the system allocates two Step 3 grants, and each user randomly selects one of them for transmission.
  • Single-root preamble sequence-based early collision detection (S-eCD) [TVT21_SeCD]: This scheme uses a single root ZC sequence and assigns a distinct prime number and cyclic shift to each user such that their product maps to a unique correlation index. Since the BS can distinguish different prime numbers regardless of cyclic shift, users selecting the same prime number are identified as collided. The scheme detects collisions at an early stage and forces collided users to back off.
@@ V-A Preamble Collision Classification Performance
This subsection demonstrates the effectiveness of the proposed early preamble collision classifier in LEO satellite communication systems. For comparison, preamble collision classifiers originally developed for terrestrial networks, namely the DRA [TWC21_DLRA] and the DCRA [TWC23_DLDouble], are adopted as benchmarks.
Table II presents the classification accuracy of the preamble collision classifiers under different channel models, antenna configurations, and values of $K$. Recall that the classifier has $(K+1)$ output classes, corresponding to 0 through $K$ users selecting the same preamble. Under the TDL-B channel model, all three classifiers exhibit similar classification accuracy. Since TDL-B is a NLoS channel model, it does not exhibit a distinct peak in the ZCZ region, which limits the effectiveness of the 1D convolutional layer. In such cases, deeper and wider neural networks, as used in DRA and DCRA, can slightly improve classification by compensating for the weak structural features. In contrast, under the TDL-D channel model, which is dominated by LoS propagation, a different trend emerges—one that is most likely in LEO SatCom scenarios. The proposed classifier consistently outperforms the others in terms of classification accuracy. This improvement can be attributed to the more prominent peak observed in the ZCZ region under LoS-dominant channels, which makes the correlation value patterns around the peak more distinguishable when a collision occurs. Moreover, the performance gap widens as the number of antennas increases. This is because DRA and DCRA average correlation values across antennas to enhance stability, whereas the proposed classifier uses all per-antenna correlation values, allowing it to exploit richer spatial information.
Table III presents the number of parameters required by each preamble collision classifier for different numbers of antennas and class settings ($K$). For DRA and DCRA, the number of parameters remains constant regardless of the number of antennas, as the input data size does not vary with antenna count. On the other hand, the proposed classifier’s parameter count increases with the number of antennas due to the expanded input size. However, while DRA and DCRA use only fully connected layers with dense inter-node connections, the proposed classifier primarily employs convolutional filters. This design drastically reduces the number of required parameters. For instance, with 8 antennas and $K=6$, the proposed classifier requires about 73 times fewer parameters than DRA and nearly 2,600 times fewer parameters than DCRA. As a result, the proposed model exhibits lower complexity and is thus more suitable for LEO satellite environments, where the onboard processing capability may be constrained [WCM24_LEOhardware]. In summary, the proposed classifier achieves comparable or even superior classification accuracy with lower computational complexity.

### v2

@@ IV-A Preamble Transmission and Reception
where $P_{d}$ is the transmit power, and $\hat{\tau}_{d}$ and $\hat{f}_{d}$ denote the estimated timing advance and Doppler shift, respectively. $T_{\mathrm{s}}$ denotes the sampling period. We consider a multi-path channel model consisting of $L_{d}$ paths for the $d$th user. The SBS is equipped with $N_{\mathrm{ant}}$ antennas, each of which receives the superposition of signals reflected through multiple paths. Each path is characterized by a channel coefficient $h_{d,\ell}^{j}$ and a propagation delay $\tau_{d,\ell}^{j}$ at the $j$th antenna. Over the short random-access time scale, we assume a quasi-static multipath channel. This is because the considered satellite channel is line-of-sight (LoS)-dominant with negligible scattering near the satellite, and the device is assumed to be stationary during an access attempt [7]. The signal received at antenna $j$ of the SBS is modeled as
@@
where $D$ denotes the total number of active users that attempt random access in the considered random access occasion, $\tau_{d}^{\mathrm{e}}$ and $f_{d}^{\mathrm{e}}$ represent the residual errors in timing and frequency pre-compensation for user $d$, and $w[n]$ is additive white Gaussian noise with zero mean and variance $\sigma_{w}^{2}$. Following the conventional detection procedure, the SBS computes the correlation between the received signal and each ZC sequence. The correlation at lag $m$ for root index $r$ at antenna $j$ is given by
@@ IV-B Preamble Collision Classifier
As mentioned, collisions are not detected by traditional schemes until Step 3. To detect preamble collisions at an early stage, we design a deep learning-based classifier that analyzes the correlation values in (8). Here, $K$ is a design parameter that specifies the maximum collision multiplicity explicitly distinguished by the classifier, and it is independent of the total number of active users $D$ in the system. For a given root index and cyclic shift, the classifier considers a total $(K+1)$ classes defined as follows:
@@
  • Class 1: Single user case, where exactly one user has used the preamble (no collision).
@@
  • Class $\boldsymbol{k}$ ($\boldsymbol{2\leq k\leq K-1}$): Exactly $k$ users have used the preamble (collision).
@@
  • Class $\boldsymbol{K}$: $K$ or more users have used the preamble (collision).
@@
The advantage of post-processing the correlation output, as opposed to relying solely on traditional threshold-based detection, is that it not only identifies whether a collision has occurred but also estimates the collision multiplicity up to the designed cut-off $K$ (with class $K$ representing $K$ or more users).
Figure 4: Representative correlation patterns within one ZCZ for a two-user preamble collision: (a) Case 1: two separated peaks (distinct arrivals), (b) Case 2: a single dominant peak due to constructive superposition, and (c) Case 3: no prominent peak due to destructive superposition.
@@
In realizing such a classifier, our goal is to propose a framework that can be easily integrated into legacy random access procedures by maintaining the core signaling of the conventional random access framework. Consequently, the input data for the classifier should be correlation values (8). Since each user independently selects a preamble and collisions across different preambles are independent events, we use a single ZCZ as the input data. A key point to note, as previously mentioned, is that in SatCom scenarios, timing advance is pre-compensated by the users during the preamble transmission stage. As shown in (5), the ZCZ size in SatCom scenarios is expected to be considerably smaller than in terrestrial networks, primarily because the otherwise large RTT uncertainty is mitigated through timing advance pre-compensation. For example, while the typical ZCZ size in terrestrial networks is approximately 24 samples [20], it is assumed to be only 8 samples in SatCom scenarios [21, 22]. This limited ZCZ size imposes a fundamental constraint on the amount of information that can be extracted for preamble collision detection. More specifically, the smaller number of samples reduces the temporal resolution available to distinguish multiple colliding signals. As a result, subtle differences in time alignment and energy distribution between users become harder to capture, making it more challenging to detect collisions based solely on the correlation values. In particular, even for the same collision multiplicity, the resulting correlation sequence within a short ZCZ can vary significantly depending on the relative arrival timing and phase among users, which may create ambiguous observations for simple rule-based detection.
To make this point intuitive, Fig. 4 illustrates representative correlation patterns within one ZCZ for a two-user collision, where two users select the same preamble. In Case 1 (Fig. 4(a)), two sufficiently separated arrivals produce two distinguishable peaks, making the collision visually apparent. In Case 2 (Fig. 4(b)), closely aligned arrivals can constructively combine into a single dominant peak, which may resemble a collision-free single-user case. In Case 3 (Fig. 4(c)), destructive superposition suppresses prominent peaks, which can even resemble an idle case. These examples show that collision-induced signatures are not necessarily captured by a single peak-based statistic and that robust collision inference requires exploiting local structures in the ZCZ.
@@
This challenge motivates leveraging additional observations beyond a single correlation trace. Fortunately, LEO satellites employ multi-antenna systems [18], which provide multiple observations of the transmitted preambles. In terrestrial networks, the correlation values obtained from multiple antennas or ports are typically averaged for threshold-based detection [14]. Similarly, existing deep learning-based schemes have also used the averaged correlation values as input data [20, 43]. Averaging enables receive antenna diversity but at the cost of reduced input dimensionality, which hinders full utilization of the original signal information. To fully exploit the information contained within preamble signals received across multiple antennas, we use the correlation values from each antenna as input data for the classifier. This preserves the spatial and temporal diversity in the received signals, providing the classifier with richer information for more accurate collision detection.
@@
However, using antenna-wise correlation values increases the input dimensionality, which requires an appropriate feature extractor to avoid overfitting and to robustly capture collision-induced signatures. More importantly, the discussion in Fig. 4 indicates that collisions manifest as localized and sometimes ambiguous structures within a short ZCZ, where the informative cues appear around a few adjacent lags rather than across the entire correlation sequence. Therefore, we employ 1D convolutional layers to explicitly learn such local patterns and their spatial consistency across antennas, while providing robustness to small timing shifts within the ZCZ. In other words, convolution is adopted not only for efficiency but also because it is well matched to the underlying signal structure of collision-induced correlation features. At the same time, compared to fully connected layers, 1D convolutional layers require significantly fewer parameters due to weight sharing, which helps keep the model lightweight despite the increased input size. This is particularly attractive for practical NTN deployments where computational and memory resources may be limited.
The structure of our proposed early preamble collision classifier is shown in Fig. 5. The input data consists of the correlation values of the preamble’s ZCZ for each antenna, denoted as $c_{r}^{j}\mathopen{}\mathclose{{\left[m}}\right]$, where $m\in\mathopen{}\mathclose{{\left[\mathopen{}\mathclose{{\left(i-1}}\right)N_{\mathrm{CS}},iN_{\mathrm{CS}}-1}}\right]$. The input layer is of size $N_{\mathrm{ant}}\times N_{\mathrm{CS}}$, where $N_{\mathrm{CS}}$ is set to 8 according to the condition in (5[21]. The kernel size of the 1D convolutional layer is set to 3, which reflects the typical delay spread of the SatCom channel as reported in [7] and captures local dependencies across adjacent samples. After the input layer, the data passes through a 1D convolutional layer with a kernel size of 3, stride 1, padding 1, and 16 channels. The rectified linear unit (ReLU), defined as $\mathrm{ReLU}(x)=\max\mathopen{}\mathclose{{\left(0,x}}\right)$, is used as the activation function. The next layer is another 1D convolutional layer with the same kernel size, stride, and padding, but with 32 channels, designed to extract high-level features of the preamble collision pattern. After passing through the activation function, the data is fed into a fully connected layer for classification. The output size of the fully connected layer is $K+1$, corresponding to the number of classes. The final output values $z_{k}$ are processed using the softmax function $\sigma(z_{k})\triangleq\frac{e^{z_{k}}}{\sum^{K}_{i=0}e^{z_{i}}}$ to compute the probability of each class. Cross-entropy is used for the loss function, i.e.,
@@ IV-C Opportunistic Transmission Scheme for Step 3
    The SBS transmits the RAR message to the users, including the preamble index, PUSCH grant, and a temporal identifier linking the RAR to the original preamble. In addition, as a lightweight 3GPP standard extension, the SBS delivers an 8-bit quantized index representing the Step 3 transmission probability (e.g., via an explicit RAR extension field or a dedicated MAC control element scheduled together with the RAR), without altering the baseline random access timeline [1, 3].
@@
    Each user refers to the RAR corresponding to its transmitted preamble index and decides whether to attempt Step 3 or perform random backoff based on the received probability index.
Based on the estimated number of collided users, we now derive the Step 3 transmission probability that maximizes the random access success probability. Notably, the following derivation does not rely on any specific traffic arrival model. It is conditioned on the estimated collision multiplicity $\hat{k}$ for a given random access occasion, and thus remains valid regardless of how the set of active users is generated. This probability allows the system to balance the trade-off between minimizing unnecessary Step 3 transmissions and maintaining a high success probability, even in the presence of collisions. Let the number of collided users in the ZCZ of a given preamble be classified as $\hat{k}$, and let the Step 3 transmission probability for this preamble be denoted as $P$. For notational simplicity, we omit the preamble index in the following derivation, without loss of generality. We define successful random access as the event in which exactly one of the users who selected the same preamble attempts Step 3, while all others perform backoff. Based on this definition, the random access success probability for the preamble is given as
@@ V Simulation Results
This section evaluates the performance of the proposed deep learning-based random access framework for LEO satellite communication systems. The tapped delay line (TDL) channel model recommended for SatCom link-level evaluation [7] is employed as the channel model. We assume that GNSS-based timing advance and frequency offset pre-compensation are performed at the user side [7]. Other random access-related parameters are summarized in Table I.
@@
The dataset for training the collision classifier consists of ZCZ correlation values generated under the TDL-B and TDL-D channel models, which are representative of non-LoS (NLoS) and LoS-dominant propagation environments, respectively. We emphasize that the classifier is designed to learn collision-induced local patterns in the correlation sequence, particularly within the ZCZ, rather than to emulate every practical propagation impairment exhaustively. Data are collected for all classes from 0 to $K$, corresponding to $K+1$ collision scenarios. To accurately simulate the reception of preambles under these scenarios, appropriate transmit power and noise configurations are required. In 3GPP-compliant systems, preamble transmission is subject to open-loop power control, where the transmit power is adjusted to achieve a target received power at the base station and thereby satisfy the detection criteria [2]. To capture this in our simulation, we configure the transmit power and noise level based on 3GPP detection requirements. Specifically, we consider the 3GPP-specified targets of misdetection probability below 1% and false alarm probability below 0.1% as a reference operating requirement [14].
TABLE I: Simulation parameter settings [21, 30, 8, 27, 5].
@@
Fig. 6 evaluates the misdetection and false alarm probabilities of the proposed classifier under various signal-to-noise ratio (SNR) conditions, where the misdetection probability is defined as the likelihood of classifying a non-zero class as class 0, and the false alarm probability is the likelihood of classifying a true class 0 as another class. To satisfy 3GPP requirements [14], the minimum required SNR is approximately 13-13 dB. Accordingly, 100,000 data samples were collected at each SNRs of {13-13, 12-12, 11-11, 10-10} dB, of which 70% were used for training and 30% for testing. Additionally, as in [20], to demonstrate the generalization performance of the trained classifier, 30,000 samples were collected at each SNRs of {9-9, 8-8, 7-7, 6-6, 5-5, 4-4} dB as a separate test set. Hyperparameters including learning rate, batch size, number of epochs, and Adam optimizer settings were tuned using $K$-fold cross-validation. The final values were set to a learning rate of $10^{-3}$, a batch size of 32, 20 epochs, and Adam optimizer parameters $(\beta_{1},\beta_{2})=(0.9,0.999)$, respectively.
@@
  • Deep learning-based random access framework for terrestrial networks (DRA) [20]: This scheme employs a deep learning-based classifier for collision detection and uses a separate neural network to extract timing advance information, which is then used to identify users and allocate distinct Step 3 grants. However, it assumes that timing advance is stationary and known to users in advance, making it unsuitable for LEO SatCom systems. Therefore, for a fair and implementable comparison under LEO SatCom conditions, we adopt an LEO-compatible adaptation of DRA as follows: when a collided preamble is detected, all users who transmitted that preamble are forced to back off (i.e., no Step 3 grant is allocated for that preamble in the current attempt).
  • Deep learning-based double contention random access (DCRA) [43]: This scheme extends DRA [20] by adopting a wider and deeper neural network for improved collision classification. When two users are classified as collided, the system allocates two Step 3 grants, and each user randomly selects one of them for transmission.
  • Single-root preamble sequence-based early collision detection (S-eCD) [44]: This scheme uses a single root ZC sequence and assigns a distinct prime number and cyclic shift to each user such that their product maps to a unique correlation index. Since the BS can distinguish different prime numbers regardless of cyclic shift, users selecting the same prime number are identified as collided. The scheme detects collisions at an early stage and forces collided users to back off.
  • Multiple threshold-based early collision detection (MT-eCD) [22]: This scheme is designed for LoS-dominant LEO SatCom channels and detects early preamble collisions by exploiting the constructive and destructive interference behavior using multiple decision thresholds. The scheme detects collisions at an early stage and forces collided users to back off.
@@ V-A Preamble Collision Classification Performance
This subsection demonstrates the effectiveness of the proposed early preamble collision classifier in LEO satellite communication systems. For comparison, preamble collision classifiers originally developed for terrestrial networks, namely the DRA [20] and the DCRA [43], are adopted as benchmarks.
@@
Table II presents the classification accuracy of the preamble collision classifiers under different channel models, antenna configurations, and values of $K$. Recall that the classifier has $(K+1)$ output classes, corresponding to 0 through $K$ users selecting the same preamble. Under the TDL-B channel model, all three classifiers exhibit similar classification accuracy. Since TDL-B is a NLoS channel model, it tends to yield less pronounced local structures in the ZCZ region, which can reduce the benefit of the 1D convolutional layer. As reflected in Table II, in such cases, deeper and wider neural networks, as used in DRA and DCRA, may slightly improve classification by compensating for the weak structural features. In contrast, under the TDL-D channel model, which is dominated by LoS propagation, a different trend emerges, which is representative of LoS-dominant settings often considered in LEO SatCom scenarios. The proposed classifier consistently outperforms the others in terms of classification accuracy. This improvement can be attributed to the more prominent peak observed in the ZCZ region under LoS-dominant channels, which makes the correlation value patterns around the peak more distinguishable when a collision occurs. Moreover, the performance gap widens as the number of antennas increases. This is because DRA and DCRA average correlation values across antennas to enhance stability, whereas the proposed classifier uses all per-antenna correlation values, allowing it to exploit richer spatial information.
Table III presents the number of parameters required by each preamble collision classifier for different numbers of antennas and class settings ($K$). For DRA and DCRA, the number of parameters remains constant regardless of the number of antennas, as the input data size does not vary with antenna count. On the other hand, the proposed classifier’s parameter count increases with the number of antennas due to the expanded input size. However, while DRA and DCRA use only fully connected layers with dense inter-node connections, the proposed classifier primarily employs convolutional filters. This design drastically reduces the number of required parameters. For instance, with 8 antennas and $K=6$, the proposed classifier requires about 73 times fewer parameters than DRA and nearly 2,600 times fewer parameters than DCRA. This reduction directly translates into lower on-board compute and memory requirements, which is particularly beneficial for practical NTN deployments subject to stringent payload constraints [41]. In summary, the proposed classifier provides a favorable accuracy–complexity trade-off, achieving comparable or even superior classification accuracy with substantially reduced model size and implementation cost.